# Combating Partial Perception Deficit in Autonomous Driving with Multimodal LLM Commonsense

### v1

@@ Abstract
Partial perception deficits can compromise autonomous vehicle safety by disrupting environmental understanding. Current protocols typically respond with immediate stops or minimal-risk maneuvers, worsening traffic flow and lacking flexibility for rare driving scenarios. In this paper, we propose LLM-RCO, a framework leveraging large language models to integrate human-like driving commonsense into autonomous systems facing perception deficits. LLM-RCO features four key modules: hazard inference, short-term motion planner, action condition verifier, and safety constraint generator. These modules interact with the dynamic driving environment, enabling proactive and context-aware control actions to override the original control policy of autonomous agents. To improve safety in such challenging conditions, we construct DriveLM-Deficit, a dataset of 53,895 video clips featuring deficits of safety-critical objects, complete with annotations for LLM-based hazard inference and motion planning fine-tuning. Extensive experiments in adverse driving conditions with the CARLA simulator demonstrate that systems equipped with LLM-RCO significantly improve driving performance, highlighting its potential for enhancing autonomous driving resilience against adverse perception deficits. Our results also show that LLMs fine-tuned with DriveLM-Deficit can enable more proactive movements instead of conservative stops in the context of perception deficits.
footnotetext: ${}^{\text{1}}$niversity at Buffalo, SUNY. ${}^{\text{2}}$niversity of Notre Dame. Correspondence to: Yuting Hu $<$yhu54@buffalo.edu$>$, Jinjun Xiong $<$jinjun@buffalo.edu$>$
Preprint.
@@ 1 Introduction
Autonomous driving has transformed the transportation landscape with its rapid evolution these years. Modern autonomous driving systems rely heavily on sensor data inputs from cameras and LiDARs, processed by deep neural networks, to enable perception, understanding, and interaction with the environment [17, 26, 7, 25]. Since perception directly influences critical driving decisions, partial information deficits caused by sensor failures or attacks can be fatal, leading to catastrophic consequences by impairing essential functions such as object detection and spatial sensing  [8, 31, 39, 46].
Figure 1: With perception deficits of safe-critical objects, the deep learning-based AV agents tend to take over-conservative or unsafe maneuvers. In contrast, LLM-RCO draws on LLM commonsense knowledge to infer possible hazards in the deficit area and plan appropriate maneuvers to solve the context.
@@
Considering severe driving scenarios with perception deficits, as illustrated in Fig. 1, the loss of safety-critical object information can undermine driving safety. Conventional safety protocols often default to immediate stopping or minimal maneuvers, which can unnecessarily impede traffic flow and still fall short when confronting partial information deficits in rare driving scenarios. Moreover, most fail-safe strategies still depend on complete perception inputs, which can be unreliable under partial perception deficits, limiting their effectiveness [45]. Human drivers demonstrate that safe navigation is possible with limited visibility by leveraging commonsense reasoning and experience to distinguish or infer critical visual information [14]. For instance, obstructions in peripheral areas may not impact immediate driving decisions, and cautious proceeding remains viable when maintaining sufficient physical distance from deficit regions or when these deficit regions show temporal stability. This human-inspired insight suggests that autonomous systems should conduct proactive context-aware operations while ensure safety. While commonsense plays a crucial role in human judgment to ensure a smooth and safe driving experience, current autonomous driving research has yet to explore how to integrate commonsense into decision-making under such challenging conditions.
@@
To address this challenge, we turn our attention to human-level intelligent multimodal large language models (LLMs), which have demonstrated remarkable reasoning capabilities and extensive commonsense knowledge across diverse applications [18, 51, 37]. Trained on vast web-scale corpora, multimodal LLMs inherently encode a wealth of traffic rules and driving knowledge, potentially equipping them with a strong driving commonsense. However, effectively leveraging this knowledge for autonomous driving remains an open question. It is not as simple as designing a prompt, particularly when driving conditions vary under different partial perception deficits. Moreover, unlike end-to-end LLM driving agents [41, 29], which can be trained through imitation learning from rule-based experts, there is no well-established expert to provide optimal strategies for vehicle movement under incomplete perception. Most fail-safe mechanisms simply default to an immediate stop, making it challenging to collect training data for an end-to-end deep learning model that directly guides autonomous vehicle movement in scenarios with perception deficits.
@@
In this paper, we present LLM-Guided Resilient Control Override (LLM-RCO), a novel framework that harnesses the commonsense knowledge and reasoning capabilities of multimodal large language models to enhance the resilience and safety of autonomous vehicle control under partial perception deficits. Specifically, when camera vision is compromised, LLM-RCO overrides autonomous vehicle control to mitigate the risk of hazardous actions resulting from erroneous predictions on compromised perception data, ensuring safer operation in adverse conditions. Inspired by human driver’s chain of thought when facing perception deficits, our framework decomposes this reasoning process into four key LLM-enhanced modules that dynamically interact with the driving environment: (1) Hazard Inference Module, which assesses potential hazards within deficit regions with past camera frames; (2) Short-Term Motion Planner, which creates a variable-length sequence of action-condition pairs adapting to the driving context and hazard inference results; (3) Action Condition Verifier, which evaluates the consistency of information deficits and immediate hazards in the current observation to confirm action feasibility, enabling safe and reliable vehicle control actions; (4) Safety Constraints Generator, which generates vehicle control limits based on safety-critical factors such as weather, lighting conditions, and traffic density; Additionally, we construct the DriveLM-Deficit dataset based on DriveLM-GVQA [42] to fine-tune the LLM for hazard inference and planning, enhancing the subsequent short-term motion planning process. To our knowledge, this paper is the first to provide an LLM-based framework for autonomous driving against perception deficits.
Contributions. We make the following contributions:
  • We propose LLM-RCO, a framework that deconstructs the human driver’s chain of thought when facing perception deficits into four LLM-fostered modules, enabling proactive vehicle movement.
@@
  • We construct DriveLM-Deficit, a dataset composed of 53, 895 driving videos with perception deficits and annotations of hazards and motion planning. We benchmark the performance of advanced LLMs on DriveLM-Deficit, highlighting their limitations.
  • We fine-tune Qwen2-VL-2B-Instruct on the DriveLM-Deficit dataset using LoRA to enhance hazard inference and motion planning of LLM-RCO.
@@
  • We validate LLM-RCO’s effectiveness in closed-loop environments using the CARLA [12] simulator separately with TransFuser [34] and InterFuser [40] agents, demonstrating that LLM-RCO consistently enhances driving scores by adopting a cautious yet proactive driving style.
Figure 2: LLM-RCO framework overview.
2 Related Works
2.1 LLM in Autonomous Driving
Recent advancements have integrated large language models (LLMs) into autonomous driving systems, offering improvements across multiple areas. LLMs have been applied as autonomous driving agents in closed-loop scenarios, demonstrating their potential for independent task handling in controlled environments [29, 30, 19, 27, 41]. They also play a critical role in enhancing autonomous driving security by detecting and mitigating adversarial attacks, making vehicles more resilient to malicious disruptions [22, 43, 10, 3, 2]. Furthermore, LLMs contribute to explainability, providing insights into decision-making processes that foster safer and more transparent vehicle control operations [9, 33, 49]. In addition, embedding LLMs into existing autonomous driving technique stacks enables commonsense, knowledge-based decision-making, enhancing autonomous vehicle’s ability to make context-aware driving decisions [11, 38, 48, 13]. Despite these advancements, there remains a significant gap in addressing motion planning under conditions of partial perception deficit like attacks or sensor failures, where LLMs could strengthen autonomous driving resilience in real-world scenarios.
2.2 Fail-Safe Strategies of Autonomous Driving
@@
Fail-safe strategies in autonomous driving rely on multi-sensor fusion [26, 6], safe degraded mode [28, 20], and emergency stop or pull-over maneuvers [5, 15] to ensure safety under perception failures and sensor malfunctions. However, these approaches have notable limitations: rule-based protocols struggle with long-tail scenarios, emergency stops disrupt traffic flow, and perception-dependent fail-safe motion planning remains vulnerable to sensor failures and adversarial attacks [28, 35, 23, 32]. While some AV frameworks use hard-coded safety control rules, such as overriding stops in low-confidence traffic light detections, these rigid mechanisms lack adaptability in real-world conditions [36, 50, 24]. A more robust approach should leverage context-aware decision-making rather than fixed rule-based overrides. This motivates LLM-RCO, which aims to provide a more effective and adaptable solution for handling perception failures in autonomous driving systems.
3 LLM-RCO
@@
In this section, we propose the LLM-RCO, which features the incorporation of LLM commonsense in multiple steps of planning and reasoning processes to enhance the resilience and safety of autonomous vehicle control strategies.
3.1 Preliminary
@@
We first define the final executed action of autonomous driving at time $t$ as $\mathcal{A}_{t}$. Action contains three vehicle control parameters: throttle, brake, and steering wheel, i.e., $\mathcal{A}_{t}=[\textit{throttle}_{t},\textit{brake}_{t},\textit{steer}_{t}]\in[0,1]\times[0,1]\times[-1,1]$. We then define an action sequence as a collection of pairs, each consisting of a tentative candidate action and its execution condition, spanning from time $t$ to time $t+n$, denoted as $\textit{S}_{t}=\{(\mathcal{C}_{t},\mathcal{A}_{t}),\cdots,(\mathcal{C}_{t+n},\mathcal{A}_{t+n})\}$. Finally, we define the environmental information at $t$ as $E_{t}=[\textit{Perception}_{t},\textit{Navi}_{t},\textit{Surrounding}_{t}]$, where $\textit{Perception}_{t}=\{I^{L}_{t},I^{F}_{t},I^{R}_{t}\}$ is the real-time images separately collected by left, front, and right camera sensors; $\textit{Navi}_{t}$ is the navigation contains the coordinate of the target point, current direction, and road geometry; and $\textit{Surrounding}_{t}$ is the surrounding environmental parameters that describe traffic, weather and daylight information. We define the safety conditions of the vehicle at time $t$ as $\textit{SC}_{t}$, describing constraints such as the maximum following distance and so on.
3.2 Framework Overview
@@
Unlike the LLM-based autonomous driving agents that employ continuous, real-time action inference, our LLM-RCO framework adopts a proactive, plan-ahead strategy. By reducing the frequency of interactions with the LLM, it lowers inference latency, leading to more efficient and smoother driving. Fig. 2 illustrates overall the LLM-RCO framework, which consists of two phases: planning and acting.
@@
LLM-RCO planning first launches a Hazard Inference Module (Section 4.1), an LLM agent that takes multiple frames of sensor context information $\textit{Percetion}_{\{t-k,\cdots,t\}}$, and then concludes the potential hazard and its movement into $\textit{Hazard}_{t}$ in Eq. (3). Then, a Short-term Motion Planning Module (Section 4.2), another LLM agent takes concluded hazard information $\textit{Hazard}_{t}$, navigation information $\textit{Navi}_{t}$ and current time perception information $\textit{Perception}_{t}$, and then generates a variable-length (we denote as $n$) condition-action sequence as $\textit{S}_{t}$.
@@
The condition-action pairs will first be verified by the Action Condition Verifier (Section 4.3), which determines whether the current environment perception $E_{t}$ is suitable for executing the condition-action $(\mathcal{C}_{t},\mathcal{A}_{t})$. If suitable, LLM-RCO executes the action $\mathcal{A}_{t}$ by the acting phase. If the Action Condition Verifier denies action $\mathcal{A}_{t}$ or the $\textit{S}_{t}$ is empty, then LLM-RCO will start a new round of planning.
4 LLM-RCO Key Components
4.1 Harzard Inference
Partial perception loss can severely impact driving when the AV system cannot detect potential hazards in compromised areas. Inferring these hazards in the perception deficit area enables the system to make more informed driving decisions. We define a hazard based on the object and its movement. Moreover, multiple hazards (denoted as $N_{t}$) may exist simultaneously. Formally:
$\textit{hazard}_{t}=\{[\textit{object}_{i},\textit{motion}_{i}]\arrowvert_{i=1}^{N_{t}}\}.$ (3)
In LLM-RCO, the hazard inference module is essential for ensuring safe driving decisions by compensating for perception loss. Errors in hazard inference can adversely affect short-term planning. To address this, we fine-tune LLMs to jointly infer potential hazards within compromised image regions and the corresponding short-term planning strategies mentioned in Section 4.2, as the pre-trained LLMs have already learned robust visual and linguistic features from large-scale datasets. Assuming that sensor failures or attack patterns remain consistent in the short term, tracking the temporal motion of the unaffected surroundings provides vital insights into their evolving dynamics, enabling the LLM to produce more accurate hazard predictions. Consequently, we input a sequence of the past $k$ multi-view images from the cameras, allowing the model to infer potential hazards and appropriate motion strategy in the deficit region based on historical context.
@@ 4.1 Harzard Inference
4.2 Short-term Motion Planing
Figure 3: We show two cases of movements using LLM-RCO under Move and Stop-Observe-Move strategies.
Short-term Motion Planner employs two distinct operational strategies $plan$ that emulate human-like adaptability: Move and Stop-Observe-Move. As demonstrated in Fig. 3, under Move strategy, the vehicle proceeds cautiously with short-term motions plan ahead, while Stop-Observe-Move requires temporary halting to gather additional information before proceeding to ensure safety in uncertain situations. Under the Move strategy, the system generates an adaptive-length sequence of planning steps. For Stop-Observe-Move strategy, when the LLM determines that an immediate stop is necessary, it specifies both the waiting duration $wait$ and move trigger conditions $\mathcal{C}_{\text{move}}$ to initiate a new round of motion planning for the vehicle’s movement. During the waiting period, only ‘stop’ actions are added into $\textit{S}_{t}$.
@@ 4.2 Short-term Motion Planing
Different from conventional defense strategies that only output driving control actions, LLM-RCO tasks the LLMs to generate an execution condition for each action to enhance driving safety and allow better adaptation to dynamic road conditions. In $\textit{S}_{t}$, each element from $(\mathcal{C}_{t},\mathcal{A}_{t})$ to $(\mathcal{C}_{t+n},\mathcal{A}_{t+n})$ is a condition-action pair, indicates that the action will be executed only when the condition is met. We define conditions as combinations of ‘consistent_deficit’ with either ‘no_immediate_hazard’ or ‘immediate_hazard’. When the observed deficits are inconsistent across frames, it implies extreme uncertainty about potential hazards arising from perception deficits. In such cases, the LLM is tasked with real-time motion inference by immediate replanning. Conversely, if deficit patterns remain consistent across frames, the LLM can utilize unaffected perception information and hazard inference results to generate proactive action sequences. As illustrated in Fig. 4, under consistent deficits, the LLM may plan more stable actions to enable smooth driving if there is no immediate hazard or opt for more aggressive maneuvers such as sudden brake in the presence of an immediate hazard.
@@
Figure 4: Illustration of inconsistently/consistently deficits across frames in various hazard cases. Left: Scenarios of Inconsistent Deficits: (a) spatial shift, (b) quantity realignment, and (c) deficit disappearance. Right: Scenarios of Consistent Deficits: (a) no immediate hazard, (b) immediate hazard due to a large deficit size, (c) immediate hazard from nearby objects in unaffected regions.
4.3 Action Condition Verifier
@@
At each timestep, LLM-RCO’s acting phase reads and executes the next action $\mathcal{A}_{t}$ in $\textit{S}_{t}$. To ensure the safe execution of plan-ahead actions, we design a rule-based action condition verifier to check $\mathcal{C}_{t}$ based on real-time perception observation. In LLM-RCO, we task LLM to generate a proactive action sequence to reduce LLM inference load. However, as the road conditions might be highly dynamic, we introduce a deficit consistency check to enable real-time action inference by motion replanning. As shown in Fig. 4, the deficit consistency is evaluated by examining the temporal quantity and spatial shifts of deficits observed across frames. We compare deficit regions across image frames in terms of both their quantity and relative spatial positions against the background. If the quantity of deficits does not align or if the spatial shift of deficit regions exceeds a predefined threshold, the condition check will fail, prompting instant motion replanning. If the consistency check passes, we assess immediate hazards by analyzing the ratio of deficit regions and detected traffic objects (e.g., cars, trucks, buses, bicycles, pedestrians, motorcycles using YOLOv11) to the camera image. This ratio, computed as the area of deficit regions and traffic object bounding boxes relative to the total image size, serves as a hazard proximity indicator. If it exceeds a predefined threshold (set to 0.05), it indicates an immediate hazard near the vehicle.
4.4 Safety Constraint Module
@@
In LLM-RCO, the Safety Constraint Module adjusts the action $\mathcal{A}_{t}$ from the Action Condition Verifier to ensure compliance with safety constraints. As safety constraints depend on various factors, for example, the maximum following distance in snowy conditions differs from that in normal conditions, we leverage the extensive commonsense knowledge of LLMs to improve driving safety. Specifically, based on the driving context including weather, daylight, traffic density, road geometry, etc., we task LLMs to generate a set of loose safety constraints $\mathcal{SC}_{t}$ to enhance vehicle control safety. As driving conditions remain stable over short time periods, we don’t need real-time inference for $\mathcal{SC}_{t}$ generation. Due to space limitations, more implementation details are provided in the Appendix A.1.
Figure 5: Driving-Oriented Fine-tuning Scheme.

### v2

# Driving Through Uncertainty: Risk-Averse Control with LLM Commonsense for Autonomous Driving under Perception Deficits

@@ Abstract
Partial perception deficits can compromise autonomous vehicle safety by disrupting environmental understanding. Existing protocols typically default to entirely risk-avoidant actions such as immediate stops, which are detrimental to navigation goals and lack flexibility for rare driving scenarios. Yet, in cases of minor risk, halting the vehicle may be unnecessary, and more adaptive responses are preferable. In this paper, we propose LLM-RCO, a risk-averse framework leveraging large language models (LLMs) to integrate human-like driving commonsense into autonomous systems facing perception deficits. LLM-RCO features four key modules interacting with the dynamic driving environment: hazard inference, short-term motion planner, action condition verifier, and safety constraint generator, enabling proactive and context-aware actions in such challenging conditions. To enhance the driving decision-making of LLMs, we construct DriveLM-Deficit, a dataset of 53,895 video clips featuring deficits of safety-critical objects, annotated for LLM fine-tuning in hazard detection and motion planning. Extensive experiments in adverse driving conditions with the CARLA simulator demonstrate that LLM-RCO promotes proactive maneuvers over purely risk-averse actions in perception deficit scenarios, underscoring its value for boosting autonomous driving resilience against perception loss challenges.
@@ 1 Introduction
Modern autonomous driving systems rely on sensor data inputs from cameras and LiDARs, processed by deep neural networks, to enable perception, understanding, and interaction with the environment (hu2023planning; liu2023real; casas2020implicit; li2022bevformer). Since perception directly influences driving decisions, partial perception deficits caused by sensor failures or attacks can be fatal, leading to catastrophic consequences (ceccarelli2022rgb; MIN2023120002; shafaei2018uncertainty; wang2020safety).
Considering driving scenarios with perception deficits, as illustrated in Figure 1, the loss of safety-critical object information can undermine the driving safety of autonomous agents. Conventional fail-safe protocols default to entirely risk-avoidant actions such as immediate stops. However, not all perception loss are catastrophic to autonomous driving safety, so triggering a fail-safe maneuver for every detected perception loss event is impractical and detrimental to the navigation goals (antonante2023task; chakraborty2024system). Moreover, risk-avoidant fail-safe strategies lack the flexibility to handle diverse driving scenarios and rely on fully restored perception inputs before resuming motion, further limiting their effectiveness (vom2020fail).
@@
Human drivers demonstrate that safe navigation is possible with limited visibility by leveraging commonsense reasoning and driving experience to infer critical visual information (fu2024drive). For example, obstructions in peripheral vision may not impact immediate driving decisions, and cautious proceeding remains viable when maintaining sufficient physical distance from deficit regions or when these deficit regions show temporal stability. This human-inspired insight suggests that autonomous systems should conduct risk-averse, proactive, and context-aware strategies rather than defaulting to overly conservative actions. While commonsense plays a crucial role in human judgment to ensure a smooth and safe driving experience, current autonomous driving research has yet to explore how to integrate commonsense into decision-making under such challenging conditions.
Figure 1: With perception deficits, deep learning-based AV agents take either over-conservative or unsafe actions, while LLM-RCO draws on LLM commonsense to infer possible hazards and plan safe motions.
@@
Multimodal large language models (LLMs) have demonstrated remarkable reasoning capabilities and extensive commonsense knowledge across diverse applications (huang2022towards; zhao2024large; sap-etal-2020-commonsense). Trained on vast web-scale corpora, multimodal LLMs inherently encode a wealth of traffic rules and driving knowledge, potentially equipping them with a strong driving commonsense. However, effectively leveraging this knowledge for autonomous driving remains an open question. Unlike end-to-end LLM driving agents (shao2024lmdrive; mao2023gpt), which can be trained through imitation learning from rule-based pilots, there is no well-established expert to provide optimal strategies for proactive movement under incomplete perception. Most fail-safe mechanisms simply default to simple stops, making it challenging to collect training data for a deep learning model that can guide autonomous vehicle movement with perception deficits.
@@
To address these challenges, we propose LLM-Guided Resilient Control Override (LLM-RCO), a risk-averse framework that harnesses the commonsense and reasoning capabilities of multimodal LLMs to enhance the resilience and safety of autonomous vehicles under partial perception deficits. LLM-RCO overrides autonomous vehicle control to mitigate the risk of hazardous actions resulting from erroneous predictions on compromised perception data, ensuring safer operation in adverse conditions. Drawing from human drivers’ reasoning in perception deficit scenarios, we break down this process into four modules in LLM-RCO: (1) Hazard Inference Module, which evaluates potential risks in deficit areas using past camera frames; (2) Short-Term Motion Planner, which generates a flexible sequence of action-condition pairs tailored to the driving context and hazard inference outcomes; (3) Action Condition Verifier, which checks the consistency of information deficits and immediate hazards in current observations to ensure action feasibility for safe and reliable vehicle control; (4) Safety Constraints Generator, which sets vehicle control limits based on safety-critical factors like weather, lighting, and traffic density. Additionally, we construct the DriveLM-Deficit dataset based on DriveLM-GVQA (sima2023drivelm) to fine-tune the LLM for hazard inference and planning, enhancing the subsequent short-term motion planning process.
Contributions.
  • We propose LLM-RCO, a risk-averse framework leveraging LLM commonsense and reasoning capabilities to enable proactive movement under perception deficits.
@@
  • We build DriveLM-Deficit, a dataset having 53, 895 driving videos with perception deficits of safety-critical objects. We benchmark the performance of advanced LLMs on DriveLM-Deficit, highlighting their limitations. We fine-tune Qwen2-VL-2B-Instruct on DriveLM-Deficit using LoRA to enhance hazard inference and motion planning of LLM-RCO.
@@
  • We validate LLM-RCO in closed-loop environments using the CARLA (dosovitskiy2017carla) simulator separately with TransFuser (prakash2021multi) and InterFuser (shao2022interfuser) agents, showing that LLM-RCO consistently enhances driving scores by adopting a cautious yet proactive driving style.
Figure 2: LLM-RCO overview.
2 Related Work
LLM in Autonomous Driving. Recent work has integrated multimodal LLMs into autonomous driving, improving agent performance in closed-loop settings (mao2023gpt; mao2023language; huang2024drivlme; long2024vlm; shao2024lmdrive), enhancing security against adversarial attacks (kong2024superalignment; song2024enhancing; chung2024towards; aldeen2024wip; aldeen2024initial), and supporting explainability in decision-making (chen2024driving; peng2024lc; xu2024drivegpt4). LLMs also enable commonsense, knowledge-driven decisions when integrated into existing stacks (10297415; sha2023languagempc; wang2024omnidrive; fang2024towards). However, motion planning under partial perception deficits, such as attacks or sensor failures, remains an underexplored area where LLMs could further enhance driving resilience.
@@ 2 Related Work
Fail-Safe Strategies in Autonomous Driving. Fail-safe strategies in autonomous driving are typically composite with multi-sensor fusion (liu2023real; cao2021invisible), degraded modes (magdici2016fail; jiang2024enhancing), and risk-avoidant emergency maneuvers such as stops or pull-overs (bogdoll2022anomaly; gao2021autonomous). However, these methods have key limitations: rule-based logic struggles with long-tail events, emergency stops disrupt traffic, and perception-dependent planning remains vulnerable to failures and attacks (magdici2016fail; ramanagopal2018failing; koopman2016challenges; pek2020fail). Hard-coded safety rules, such as overriding stops for uncertain traffic light detections, lack adaptability in dynamic conditions (ren2019security; yurtsever2020survey; koopman2017autonomous). This motivates LLM-RCO, which aims to provide a more risk-averse and flexible solution for handling perception failures in autonomous driving systems.
3 Risk-averse Control: LLM-RCO
@@
In this section, we propose LLM-RCO, which features the incorporation of LLM commonsense in multiple steps of planning and reasoning processes to enhance the resilience and safety of autonomous vehicle control strategies.
@@
Preliminary. We define the action of autonomous driving at time $t$ as $\mathcal{A}_{t}$. Action contains three vehicle control parameters: throttle, brake, and steering wheel, i.e., $\mathcal{A}_{t}=[\textit{throttle}_{t},\textit{brake}_{t},\textit{steer}_{t}]\in[0,1]\times[0,1]\times[-1,1]$. An action sequence is a collection of pairs, each consisting of a tentative candidate action and its execution condition, spanning from time $t$ to time $t+n$, denoted as $\textit{S}_{t}=\{(\mathcal{C}_{t},\mathcal{A}_{t}),\cdots,(\mathcal{C}_{t+n},\mathcal{A}_{t+n})\}$. The environmental information at $t$ is $E_{t}=[\textit{Perception}_{t},\textit{Navi}_{t},\textit{Surrounding}_{t}]$, where $\textit{Perception}_{t}=\{I^{L}_{t},I^{F}_{t},I^{R}_{t}\}$ is the real-time images separately collected by left, front, and right camera sensors; $\textit{Navi}_{t}$ is the navigation contains the coordinate of the target point, current direction, and road geometry; and $\textit{Surrounding}_{t}$ is the surrounding environmental information that describe traffic, weather, and daylight information. The safety conditions of the vehicle at time $t$ is $\textit{SC}_{t}$, describing constraints such as the maximum following distance.
@@
Framework Overview. Unlike the LLM-based autonomous driving agents that employ continuous, real-time action inference, our LLM-RCO framework adopts a proactive, plan-ahead strategy. By reducing the frequency of interactions with the LLM, it lowers inference latency, leading to more efficient and smoother driving. Figure 2 shows the LLM-RCO framework, which consists of planning and acting phases.
@@
LLM-RCO planning first launches a Hazard Inference Module, an LLM agent that takes multiple frames of sensor context information $\textit{Percetion}_{\{t-k,\cdots,t\}}$, and concludes the potential hazard and its movement into $\textit{Hazard}_{t}$ in Eq. equation 3. Then, a Short-term Motion Planning Module, another LLM agent takes concluded hazard information $\textit{Hazard}_{t}$, navigation information $\textit{Navi}_{t}$ and current time perception information $\textit{Perception}_{t}$, and then generates a variable-length (we denote as $n$) condition-action sequence as $\textit{S}_{t}$.
@@
The condition-action pairs will first be verified by the Action Condition Verifier, which determines whether the current environment perception $E_{t}$ is suitable for executing the condition-action $(\mathcal{C}_{t},\mathcal{A}_{t})$. If suitable, LLM-RCO executes the action $\mathcal{A}_{t}$ by the acting phase. If the Action Condition Verifier denies action $\mathcal{A}_{t}$ or the $\textit{S}_{t}$ is empty, then LLM-RCO will start a new round of planning.
4 Key Components in LLM-RCO
Hazard Inference. Partial perception loss can severely impact driving when the AV system cannot detect potential hazards in compromised areas. Inferring these hazards in the perception deficit area enables the system to make more informed driving decisions. We define a hazard based on the object and its movement. Moreover, multiple hazards (denoted as $N_{t}$) may exist simultaneously. Formally:
[…]
In LLM-RCO, the hazard inference module is essential for ensuring safe driving decisions by compensating for perception loss. To enhance the hazard inference accuracy, we fine-tune LLMs to jointly infer potential hazards within compromised image regions and the corresponding short-term planning strategies. Assuming that sensor failures or attack patterns remain consistent in the short term, tracking the temporal motion of the unaffected surroundings provides vital insights into their evolving dynamics, enabling the LLM to produce more accurate hazard predictions. Consequently, we input a sequence of the past $k$ multi-view images from the cameras, allowing the model to infer potential hazards and plan with historical context.
@@ 4 Key Components in LLM-RCO
Figure 3: Driving visualization with LLM-RCO under Move and Stop-Observe-Move strategies.
Short-term Motion Planing. Short-term Motion Planner employs two distinct operational strategies planplan that emulate human-like adaptability: Move and Stop-Observe-Move. As demonstrated in Figure 3, under Move strategy, the vehicle proceeds cautiously with short-term motions plan ahead, while Stop-Observe-Move requires temporary halting to gather additional information before proceeding to ensure safety in uncertain situations. Under the Move strategy, the system generates an adaptive-length sequence of planning steps. For Stop-Observe-Move strategy, when the LLM determines that an immediate stop is necessary, it specifies both the waiting duration waitwait and move trigger conditions $\mathcal{C}_{\text{move}}$ to initiate a new round of motion planning for the vehicle’s movement. During the waiting period, only ‘stop’ actions are added into $\textit{S}_{t}$.
@@
Different from conventional fail-safe strategies that only output control actions, LLM-RCO tasks the LLMs to generate an execution condition for each action to enhance driving safety and allow better adaptation to dynamic road conditions. In $\textit{S}_{t}$, each element from $(\mathcal{C}_{t},\mathcal{A}_{t})$ to $(\mathcal{C}_{t+n},\mathcal{A}_{t+n})$ is a condition-action pair, indicates that the action will be executed only when the condition is met. We define conditions as combinations of ‘consistent_deficit’ with either ‘no_immediate_hazard’ or ‘immediate_hazard’. When the observed deficits are inconsistent across frames, it implies extreme uncertainty about potential hazards arising from perception deficits. In such cases, the LLM is tasked with real-time motion inference by immediate replanning. Conversely, if deficit patterns remain consistent across frames, the LLM can utilize unaffected perception information and hazard inference results to generate proactive action sequences. As illustrated in Figure 4, under consistent deficits, the LLM may plan more stable actions to enable smooth driving if there is no immediate hazard or opt for more aggressive maneuvers such as sudden brake in the presence of an immediate hazard.
@@
Figure 4: Deficits across frames. Left: Scenarios of Inconsistent Deficits: (a) spatial shift, (b) quantity realignment, and (c) deficit disappearance. Right: Scenarios of Consistent Deficits: (a) no immediate hazard, (b) immediate hazard due to a large deficit size, (c) immediate hazard from nearby objects in unaffected regions.
@@
Action Condition Verifier. At each timestep, LLM-RCO’s acting phase reads and executes the next action $\mathcal{A}_{t}$ in $\textit{S}_{t}$. To ensure the safe execution of plan-ahead actions, we design a rule-based action condition verifier to check $\mathcal{C}_{t}$ based on real-time perception observation. As the road conditions might be highly dynamic, we introduce a deficit consistency check to enable real-time action inference by motion replanning. As shown in Figure 4, the deficit consistency is evaluated by examining the temporal quantity and spatial shifts of deficits observed across frames. We compare deficit regions across image frames in terms of both their quantity and relative spatial positions against the background. If the quantity of deficits does not align or if the spatial shift of deficit regions exceeds a predefined threshold, the condition check will fail, prompting instant motion replanning. If the consistency check passes, we assess immediate hazards by analyzing the ratio of deficit regions and detected traffic objects (e.g., cars, trucks, buses, bicycles, pedestrians, motorcycles using YOLOv11) to the camera image. This ratio, computed as the area of deficit regions and traffic object bounding boxes relative to the image size, serves as a hazard proximity indicator. If it exceeds a predefined threshold (set to 0.05), it indicates an immediate hazard near the vehicle.
@@
Safety Constraint Module. In LLM-RCO, the Safety Constraint Module adjusts the action $\mathcal{A}_{t}$ from the Action Condition Verifier to ensure compliance with safety constraints. As safety constraints depend on various factors, for example, the maximum following distance in snowy conditions differs from that in normal conditions, we leverage the extensive commonsense knowledge of LLMs to improve driving safety. Specifically, based on the driving context including weather, daylight, traffic density, road geometry, etc., we task LLMs to generate a set of loose safety constraints $\mathcal{SC}_{t}$ to enhance vehicle control safety. As driving conditions remain stable over short time periods, we don’t need real-time inference for $\mathcal{SC}_{t}$ generation. Due to space limitations, more implementation details are provided in the Appendix A.2.